1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,
La lupfantomo  Le loup-garou  The Banshee (An Approach)
translated by Bertilo Wennergren  translated by R. Platteau  translated by Max Knight

Jen lupfantom’ de famili’ meznokte iam kuris for al tombo de la profesor’. kaj petis pri konjugaci’.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle ! ”

One night, a banshee slunk away

from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

Do diktis tiu dokta ul’ starante sur la kructabul’ dum lup’ senmova kiel splinto atentis vortojn de l’ mortinto.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle ! ”

The village teacher climbed up straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

De “lupfantomo”, simpla rekta, tra “lupfintomo”, plej perfekta, plu “lupfontomo”, iom rara, ĝis “lupfuntomo”, ho, koŝmara.

“ - On peut à ton sujet puisque nous y sommes Se demander encor bien des choses en somme. Loup-garoù ? mais avant Il faut se dire : loup-garquand ?

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

Ĉi participoj plaĉis multe al la fantomo, kaj rezulte ĝi petis: “Legu plu el PIV-o pri participoj en pasivo.”

Loup-garcomment ? et puis Bien évidemment loup-garqui ? ” Toutes ces questions semblèrent au loup bien absconses, Car en fait ce qu’il voulait c’était les réponses.

The banshee marveled at the cases

and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singular!”

La viro ĉagreniĝis tamen: “Jam fermis mi per fin’ kaj amen, ĉar netransiraj bestoj feblas kaj la pasiv’ neniam eblas.”

“ - Point de réponses, dit le puits de savoir, Etre imaginaire, nul ne peut te voir ! ” A ces mots l’animal éclata tout en larmes. “ - Oh!, même les baisers de ma douce moitié The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

…  …  …

2/2

Sed lupfantomo pri transiro ja spertas pli ol eĉ vampiro. La lup’ revenis larmokula al sia famili’ ulula.

Ne pourrons plus me faire me sentir tout entier. Rester un loup-garou est un sort qui m’alarme. ” Mais comme le loup était bien poli Il remercia et puis repartit.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Esperanto by Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu la retejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into French by R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

http://www.poezio.net/version?poem-id=825&version-id=1653,1655,1660